OpenStudy (anonymous):

Ratio of boysw to girls of 460 students is 11:12. Find the number of boys. Pretty sad, but how do i find this again?

OpenStudy (ganpat):

let boys be x.. so girls = 460 - x.. the ratio says, x/ (460-x) = 11/12... can u get x ??

OpenStudy (anonymous):

is that x divided by (460-x) =11/12?

OpenStudy (ganpat):

yes...

OpenStudy (ganpat):

so whats x ??

OpenStudy (anonymous):

well i did something different.... i did 11/23 x 460 and got 220. but i dont know how to do it your way...

OpenStudy (anonymous):

and your way is the way i need to do it!

OpenStudy (ganpat):

it says ratio of boys to girls of 460 students.. so x/ (460 - x) = 11/12.. so 12x = 460 (11) - 11x 12x + 11x = 460(11) 23x = 460 (11) x = 220.. = boys and girls = 460 - 220= 240... = girls..

OpenStudy (amistre64):

if the ratio is 11:12, that tells me that there are groups of (11+12) N(11+12) = 460 tells us how many groups there are N(11:12) tells us how many boys and girls there are

OpenStudy (ganpat):

@maherman97 : still confused ??

OpenStudy (anonymous):

well i dont understand where the (11) comes from...

OpenStudy (anonymous):

like you have 12x+11x=460 i understand that... but where the other (11) comes from

OpenStudy (ganpat):

no.. y 12x + 11x ??

OpenStudy (ganpat):

it says, total 460 students.. if b + g = 460 so if boys would be x, than girls should be (460 - x) .. right ??

OpenStudy (ganpat):

?????

OpenStudy (anonymous):

yeah sorry haha my bad.

OpenStudy (ganpat):

dont be sorry !!!

OpenStudy (anonymous):

thanks tho! ive got another one.. State the ratio in simplest form: three days to three weeks. would that be 3:21 which would simplify to 1:7?

OpenStudy (ganpat):

yup.. 3/ (3 * 7) .. 7 days a week.. 3/21.. = 1/7...

OpenStudy (anonymous):

okay thanks!

OpenStudy (ganpat):

W/c...